which equation represents the line that passes through the point (-6,2) and has a slope of -1?
Answer by jim_thompson5910(35256)


Answer in point-slope form: y - 2 = -1(x + 6)
Answer in slope-intercept form: y = -x - 4
Answer in standard form: x+y = -4

These are three different ways to express the same thing.
==========================================================

Work Shown:

Plug m = -1 and (x1,y1) = (-6,2) into the point slope formula. Solve for y.
y - y1 = m(x - x1)
y - 2 = -1(x - (-6))
y - 2 = -1(x + 6)
y - 2 = -1*x - 1*6 ... distribute
y - 2 = -x - 6
y - 2 + 2 = -x - 6 + 2 ... add 2 to both sides
y = -x - 4

Add x to both sides to get all the variable terms to one side. This will get us x+y = -4, which is the equation in standard form.